Is the sequence geometric? If so, identify the common ratio.
6, 12, 24, 48, ...

Answer by math_tutor2020(3817)


Answer: Yes it is geometric. The common ratio is 2.

We can determine this by dividing each term by its previous term.
r = common ratio
r = term2/term1 = 12/6 = 2
r = term3/term2 = 24/12 = 2
r = term4/term3 = 48/24 = 2
We get the same result each time, which is the common ratio.
It means we double each term to get the next term.